OpenStudy (anonymous):

Solve the system of equations. 2x+3y=13 -6x+4y=0 A. (2, 3) B. (1/2,4) C. (5, 1) D. (–2, –3)

OpenStudy (anonymous):

Do you wanna solve it by Substitution, or Elimination? ;o

OpenStudy (anonymous):

Elimination.

OpenStudy (anonymous):

Okay, well you can multiply the top equation by -3 if you would like to solve for Y first.?

OpenStudy (anonymous):

I you were solving for why first.. wouldnt you multiply 4 in the first equation?? There isnt a -3.

OpenStudy (anonymous):

Nope, Multiply by 3, because you want the First equations x Values to be 6, so the X's can cancel out. i dont know why i said -3 0-0

OpenStudy (anonymous):

I did that. 3(2x+3y=13) i got 6x+9y=39 -6x+4y=0 then i did elimination which was 0x+5y=39 5 5 and i got 7.8 O.o

OpenStudy (anonymous):

Urm, you have to add, and 9+4= 13, which will give you 13y=39. and y=3 c:

OpenStudy (anonymous):

You basically already have your answer, you really dont have to keep going if you look at your answer choices, A is the only one that has a Y value of 3 :3

OpenStudy (anonymous):

Alright i got it thanks:)

OpenStudy (anonymous):

Your welcome! (:

OpenStudy (anonymous):

Can you help me with another one @MarrBear ??

OpenStudy (anonymous):

Okay Gurl. xD

OpenStudy (anonymous):

Solve the system of equations. 8x – 3y = 1 –2x + 3y = 11 A. (–1, –3) B. (–1, 3) C. (2, 5) D. (5, 2)

OpenStudy (anonymous):

OO. Well, you have to add, which cancels out your Y's.

OpenStudy (anonymous):

youll get 6x=12. and then you will have x=2.

OpenStudy (anonymous):

Which seems to be C. :)

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

Anytime (: